# Avoiding Catastrophe in Continuous Spaces by Asking for Help

### v2

@@ Abstract
Most reinforcement learning algorithms with formal regret guarantees assume all mistakes are reversible and essentially rely on trying all possible behaviors. This approach leads to poor outcomes when some mistakes are irreparable or even catastrophic. We propose a variant of the contextual bandit problem where the goal is to minimize the chance of catastrophe. Specifically, we assume that the payoff each round represents the chance of avoiding catastrophe that round, and try to maximize the product of payoffs (the overall chance of avoiding catastrophe). We allow a limited number of queries to a mentor and assume a Lipschitz continuous payoff function. We first show that in general, any algorithm either constantly queries the mentor or is nearly guaranteed to cause catastrophe. However, when the mentor policy class has bounded Natarajan dimension and contains at least some “reasonable” policies, we provide an algorithm whose regret and rate of querying the mentor both approach 0 as the time horizon grows. We also present an alternative algorithm which provides the same regret and query guarantees when the mentor’s action changes a constant number of times in a 1D state space, and can handle adversarially chosen states.
@@ 1 Introduction
In reinforcement learning (RL), an agent finds itself in an unknown environment and wants to maximize some payoff function, commonly modeled as a Markov Decision Process (MDP). Typically, the only way for the agent to learn about its environment is by trying actions and observing their effects. This presents a challenge when some actions are inherently dangerous: for example, the agent must learn to avoid self-destructive actions without actually trying those actions.
This problem is hopeless without additional assumptions, since any untried action could lead to either paradise or disaster, and the agent has no way to predict which. Most RL theory evades this issue by either assuming ergodicity, meaning that any state is reachable from any other (e.g., [15]), or by resetting the agent at the start of each “episode” (e.g., [1]). In the real world, however, many actions lead to unrecoverable situations: we call these catastrophes. Examples could include ingesting poison or crashing a robotic vehicle. Resetting the agent may be expensive (if it destroys itself), inadequate (if it irreparably harms others), or impossible (if it is outside of human supervision).
Rather than ignore the possibility of catastrophe, we explore whether access to a mentor might enable an agent to navigate risky (i.e., non-ergodic single-episode) environments. We assume the mentor illustrates the optimal action in the current state, and we would like an algorithm which obtains near-optimal payoff with a limited number of mentor queries. “Near-optimal payoff” is often modeled via regret, defined as the gap between the optimal total payoff and the agent’s actual total payoff. For undiscounted MDPs (i.e., each time step has the same weight in the objective), one typically aims for regret sublinear in the time horizon $T$, or equivalently, for the average regret per time step to approach 0 for large $T$. Similarly, a sublinear number of mentor queries implies that the rate of querying the mentor goes to 0 for large $T$—that is, the agent gradually becomes self-sufficient. Formally, we pose the following open problem:
Is there an algorithm for non-ergodic single-episode undiscounted MDPs which ensures that both the regret and the number of mentor queries are sublinear in $T$?
An additional assumption is necessary, as the agent needs some way to make inferences about unseen states in order to decide when to query. Much past work has used Bayesian inference, which suffers tractability issues in complex environments. [^1] For the curious reader, [3] provides a thorough treatment. See also Section 1.3 for further discussion. In this paper, we use Lipschitz continuity instead, which captures the intuition that an action should have similar effects in similar states. Importantly, the Lipschitz assumption is compatible with any state space and action space with a distance metric.
@@ 1.1 Our model
In this paper, we target a more modest goal: can the agent simply avoid catastrophe? This is clearly necessary (but not sufficient) for sublinear regret in general MDPs. Furthermore, avoiding catastrophe is the precise ability which standard RL algorithms lack.
Focusing on catastrophe also allows us to streamline the model. If we redefine the payoff at time $t$ to be the probability of avoiding catastrophe on that time step, then the product of payoffs represents the overall chance of avoiding catastrophe. Representing catastrophe in the objective function eliminates the need to model non-ergodicity through the transition function, as MDPs do. Our multiplicative objective captures the same stringency: a single mistake can ruin the entire endeavor. [^2] An additive objective with unbounded payoffs could also capture the core challenge of needing to avoid catastrophic actions without ever trying them. However, we feel that the multiplicative objective is more intuitive due its interpretation as the probability of avoiding catastrophe.
@@
The resulting model is similar to the contextual bandit problem: on each time step, the agent observes a state (we generally assume i.i.d. states), selects an action (or queries the mentor), and obtains some payoff. To our knowledge, multiplicative objective functions have not been studied for contextual bandits. Lastly, unlike the standard contextual bandit model, we assume that the agent does not observe payoffs. This is because the payoff in our model represents the chance of avoiding catastrophe that round, and in the real world, one only observes whether catastrophe occurred, not its probability. [^3] One may be able to detect “close calls” in some cases, but observing the precise probability seems unrealistic.
@@ 1.2 Our results
We first show that in general, any algorithm with sublinear queries to the mentor has arbitrarily poor regret in the worst-case (Theorem 3.1). As a corollary, even when it is possible to avoid catastrophe with certainty, any algorithm either asks for help on a constant fraction of time steps or is nearly guaranteed to cause catastrophe. This holds even for randomized algorithms and simple environments: our construction uses just two actions and assumes states are uniformly random across $[0,1]$.
On the other hand, the class of mentor policies in our negative result is quite complex in the sense of having unbounded VC dimension [31]. What if we restrict the mentor policy to be relatively simple?
We present an intuitive algorithm whose regret and rate of querying the mentor both go to 0 as $T\to\infty$ when the mentor policy class has bounded VC dimension (or Natarajan dimension, more generally [^4] Natarajan dimension [24] generalizes VC dimension (which only handles binary prediction) to multi-class prediction. In our case, each class is an action, and a policy is a multi-class predictor from states to actions.) and contains at least some “reasonable” policies (Theorem 4.5). Our algorithm can handle an arbitrary unknown distribution over states, a multi-dimensional unbounded state space, and many actions. It does not need detailed access to the feature embedding (instead relying on two simple operations) and does not need to know the Lipschitz constant. It does need to know the policy class and (an upper bound on) the Natarajan dimension (as is standard in PAC learning). This is our primary positive result.
Finally, we give another algorithm with the same guarantees of sublinear queries and near-zero regret under a different simplicity assumption (Theorem 5.2). Specifically, we assume that the mentor’s action changes a constant number of times in a 1D state space. This algorithm can also handle adversarially chosen states.
We generally assume that the mentor’s policy is optimal, but our positive results can be generalized to a suboptimal mentor under a slightly different Lipschitz assumption (Remark B.1).
1.3 Related work
See [9; 11] for full surveys of safe RL. Here, we discuss only the work most relevant to our own.
@@ 1.3 Related work
Despite the ubiquity of irreversible costs in the real world, there remains a dearth of theoretical work on this topic. This may be due to the fundamental modeling question of how the agent should learn about novel states or actions without actually trying them.
The most common approach is to allow the agent to ask for help. This alone is insufficient, however: the agent must have some way to decide when to ask for help. A popular solution is to perform Bayesian inference on the world model, but this has two tricky requirements: (1) a prior distribution which contains the true world model (or an approximation), and (2) an environment where computing (or approximating) the posterior is tractable. A finite set of possible environments satisfies both conditions, but is unrealistic in many real-world scenarios. In contrast, our algorithm can handle a continuous unbounded state space, which is crucial for many real-world scenarios in which one never sees the exact same state twice.
Bayesian inference combined with asking for help is studied by [5; 6; 17; 21]. We also mention [12; 22; 30], which utilize Bayesian inference in the context of safe RL without asking for help (and without regret bounds).
We are only aware of two papers which theoretically address irreversibility without Bayesian inference: [10] and [20]. The former proposes to sample trajectories and learn reversibility based on temporal consistency between states: intuitively, if $s_{1}$ always precedes $s_{2}$, we can infer that $s_{1}$ is unreachable from $s_{2}$. Although the paper theoretically grounds this intuition, there is no formal regret guarantee. The latter presents an algorithm which asks for help in the form of rollouts from the current state. However, the regret bound and number of rollouts are both linear in the worst case, due to the dependence on the $\gamma^{*}$ parameter which roughly captures how bad an irreversible action can be. In contrast, our algorithms achieve near-zero regret even when actions can be maximally bad.
@@
Contextual bandits.
Like MDPs, the contextual bandit model involves an agent who selects an action on each time step based on the current state (or “context”) to maximize payoff. Unlike MDPs, the state on each time step is determined randomly or by an adversary, instead of by a transition function. See [29] for an overview.
The contextual bandit literature often uses Lipschitz continuity of payoff functions as a way to generalize between states, which affirms our use of that same assumption [4; 14; 18; 19; 26; 28]. However, to our knowledge, all prior contextual bandit work aims to maximize the sum of payoffs, which implicitly assumes that no mistake is catastrophic: no matter how poorly you perform, you can always make it up given a sufficiently long time horizon.
@@
Perhaps the best-studied example of a multiplicative objective is the Nash welfare function (also known as the Nash bargaining solution), defined as the geometric mean of agent utilities [16; 23]. Inspired by Nash welfare, [2] recently studied algorithms for maximizing the geometric mean of payoffs across time steps in the classic (non-contextual) multi-armed bandit problem. However, the geometric mean of payoffs is much more forgiving of errors compared to the product of payoffs. Imagine that the agent obtains a payoff of $1-1/\sqrt{T}$ per time step: then $\lim_{T\to\infty}(1-1/\sqrt{T})^{T}=0$ but $\lim_{T\to\infty}((1-1/\sqrt{T})^{T})^{1/T}=\lim_{T\to\infty}(1-1/\sqrt{T})=1$. In our setting, it is crucial that the cumulative chance of catastrophe (the product of payoffs) is small, not just the average chance of catastrophe per time step (the geometric mean of payoffs).
Multiplicative objectives also arise in the context of compound growth or repeated gambles, where one’s wealth at time $t$ can be modeled as the initial wealth times a product of growth factors (see, e.g., [25]). Our problem could be viewed as a compound growth problem where one’s “wealth” (the total probability of avoiding catastrophe) never increases. However, we are not aware of any mathematical tools related to compound growth that handle the contextual bandit aspects of our problem.
The rest of the paper is structured as follows. Section 2 formally defines our model. Section 3 presents our negative result for general mentor policies. Section 4 presents our positive result for policy classes with bounded Natarajan dimension. Section 5 presents our second positive result under a different simplicity assumption. Section 6 discusses limitations. All proofs are deferred to the appendix.
@@ 2 Model
Lipschitz contextual bandits.
Let $\mathbb{N}$ refer to the strictly positive integers and let $T\in\mathbb{N}$ be the time horizon. Let $S\subseteq\mathbb{R}^{n}$ be the state space, and let $\mathbf{s}=(s_{1},s_{2},\dots s_{T})\in S^{T}$ be a sequence of states. (The contextual bandit literature typically uses “context” instead of “state”, but we have chosen to use “state” to align with MDP terminology.) With the exception of Section 5, we assume that each state is sampled i.i.d. from an unknown distribution $\mathcal{D}$ over $S$.
For $k\in\mathbb{N}$, let $[k]=\{1,2,\dots,k\}$. On each time step $t\in[T]$, the agent must select an action $a_{t}\in A\cup\{\hat{a}\}$, which generates payoff $\mu(s_{t},a_{t})\in[0,1]$. We allow the agent’s choice of action to be randomized. The action $\hat{a}$ constitutes querying the mentor, which we will define shortly. The agent never observes payoffs: the only way to learn about $\mu$ is by querying the mentor. We assume that $|A|\leq T$, since the agent can never observe more than $T$ distinct actions.
We assume that $\mu$ is Lipschitz continuous with respect to $S$: there exists $L>0$ such $\forall s,s^{\prime}\in S,a\in A$, $|\mu(s,a)-\mu(s^{\prime},a)|\leq L||s-s^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance.
@@
The special action $\hat{a}$ constitutes querying a mentor who knows an optimal policy $\pi^{*}:S\to A$. When action $\hat{a}$ is chosen, the mentor informs the agent of the action $\pi^{*}(s_{t})$ and the agent obtains optimal payoff $\mu(s_{t},\pi^{*}(s_{t}))=\mu^{*}(s_{t})=\max_{a\in A}\mu(s_{t},a)$. Let $Q_{t}=\{i\in[t]:a_{i}=\hat{a}\}$ be the set of time steps up to and including $t$ which involve a query to the mentor.
We would like an algorithm which becomes “self-sufficient” over time: the rate of querying the mentor should go to 0 as $T\to\infty$. Intuitively, this is equivalent to the cumulative number of queries being sublinear in $T$: $|Q_{T}|\in o(T)$. However, interpreted literally, $|Q_{T}|\in o(T)$ is not a rigorous definition. This is because $|Q_{T}|$ is not just a function of $T$: it can also depend on $\mathbf{s}$, $\mu$ and $\pi^{*}$.
To formalize this, let $|Q_{T}(\mathbf{s},\mu,\pi^{*})|$ be the number of queries to the mentor up to time $T$, given payoff function $\mu$ and state sequence $\mathbf{s}$. Then we say that $|Q_{T}|$ is sublinear in $T$, denoted $|Q_{T}|\in o(T)$, if there exists $g:\mathbb{N}\to\mathbb{N}$ such that $g(T)\in o(T)$ and $|Q_{T}(\mathbf{s},\mu,\pi^{*})|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{*}$.
@@
If $\mu(s_{t},a_{t})\in[0,1]$ is the chance of avoiding catastrophe on time step $t$ (conditioned on no prior catastrophe), then $\prod_{t=1}^{T}\mu(s_{t},a_{t})$ represents the agent’s overall chance of avoiding catastrophe. [^5] Conditioning on no prior catastrophe means we do not need to assume that these probabilities are independent (and if catastrophe has already occurred, this time step does not matter). This is due to the chain rule of probability. For a time horizon $T$, payoff function $\mu$, mentor policy $\pi^{*}$, and state sequence $\mathbf{s}$, the agent’s regret is
$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
where $\mathbf{a}=\{a_{1},\dots,a_{T}\}$ is the set of actions selected by the algorithm (which depends on $\pi^{*}$). We focus on a worst-case choice of $\mu$ and $\pi^{*}$. [^6] The choice of $\pi^{*}$ can actually matter: suppose $a$ is optimal everywhere but $a^{\prime}$ is optimal only for this state.
VC and Natarajan dimension.
VC dimension [31] and Natarajan dimension [24] are standard measures of learning difficulty. The former applies to binary classes, while the latter handles multi-class prediction (and reduces to the former in the case of two classes). In our setting, a policy can be viewed as a multi-class predictor, where each action is a class. The precise definitions of these dimensions are not relevant to our paper, as we only invoke Natarajan dimension indirectly through the fundamental theorem of multi-class PAC learning (Theorem 4.4). We direct the reader to [27] for a comprehensive overview.

### v3

# Avoiding Catastrophe in Online Learning by Asking for Help

@@ Abstract
Most learning algorithms with formal regret guarantees assume that no mistake is irreparable and essentially rely on trying all possible behaviors. This approach is problematic when some mistakes are catastrophic, i.e., irreparable. We propose an online learning problem where the goal is to minimize the chance of catastrophe. Specifically, we assume that the payoff in each round represents the chance of avoiding catastrophe that round and aim to maximize the product of payoffs (the overall chance of avoiding catastrophe) while allowing a limited number of queries to a mentor. We first show that in general, any algorithm either constantly queries the mentor or is nearly guaranteed to cause catastrophe. However, in settings where the mentor policy class is learnable in the standard online learning model, we provide an algorithm whose regret and rate of querying the mentor both approach 0 as the time horizon grows. Conceptually, if a policy class is learnable in the absence of catastrophic risk, it is learnable in the presence of catastrophic risk if the agent can ask for help.
@@ 1 Introduction
There has been mounting concern over catastrophic risk from AI, including but not limited to autonomous weapon accidents [1], bioterrorism [30], and cyberattacks on critical infrastructure [14]. See [10] and [20] for taxonomies of societal-scale AI risks. In this paper, we use “catastrophe” to refer to any kind of irreparable harm. This definition also covers smaller-scale (yet still unacceptable) incidents such as serious medical errors [11], crashing a robotic vehicle [24], or discriminatory sentencing [39].
The gravity of these risks contrasts starkly with the dearth of theoretical understanding of how to avoid them. Nearly all of learning theory explicitly or implicitly assumes that no single mistake is too costly. We focus on online learning, where an agent repeatedly interacts with an unknown environment and uses its observations to gradually improve its performance. Most online learning algorithms essentially try all possible behaviors and see what works well. We do not want autonomous weapons or surgical robots to try all possible behaviors.
More precisely, trial-and-error-style algorithms only work when catastrophe is assumed to be impossible. This assumption manifests differently in different subtypes of online learning. In the standard online learning model, the agent’s actions have no permanent effect on the environment. [^1] More precisely, the state can depend on the agent’s previous actions, but the agent’s performance is always evaluated with respect to the optimal policy on the same sequence of states. Online reinforcement learning allows the agent’s actions to permanently affect the environment, but typically assumes that either no action has irreversible effects (e.g., [21]) or that the agent is reset at the start of each “episode” (e.g., [2]). One could train an agent entirely in a controlled lab setting where the above assumptions do hold, but we argue that sufficiently general agents will inevitably encounter novel scenarios when deployed in the real world. Machine learning models often behave unpredictably in unfamiliar environments (see, e.g., [32]), and we do not want AI biologists or robotic vehicles to behave unpredictably.
The goal of this paper is to understand the conditions under which it is possible to formally guarantee avoidance of catastrophe in online learning. Certainly some conditions are necessary, because if the agent can only learn by trying actions directly, the problem is hopeless: any untried action could lead to paradise or disaster and the agent has no way to predict which. In the real world, however, one needn’t learn through pure trial-and-error: one can also ask for help.
@@ 1.1 Our model
We propose an online learning model of avoiding catastrophe with mentor help. On each time step, the agent observes a state, selects an action (or queries the mentor), and obtains a payoff. [^2] Depending on the field, states are also known as inputs, instances, or contexts, actions are also known as outputs, labels, or arms, and payoffs are also known as rewards or losses. Each payoff represents the probability of avoiding catastrophe on that time step, and the agent’s goal is to maximize the product of payoffs, which corresponds to minimizing the total chance of catastrophe. We consider the product of payoffs obtained while learning, not the product of payoffs of some final policy: finding a catastrophe-avoiding policy is futile if catastrophe occurs while learning.
The (possibly suboptimal) mentor has a fixed policy, and when queried, the mentor illustrates their policy’s action in the current state. We desire an agent whose regret – defined as the gap between the mentor’s performance and the agent’s performance – approaches zero as the time horizon $T$ grows. In other words, with enough time, the agent should avoid catastrophe nearly as well as the mentor. We also expect the agent to become self-sufficient over time: formally, the number of queries to the mentor should be sublinear in $T$, or equivalently, the rate of querying the mentor should go to zero.
1.2 Our assumptions
The agent needs some way to make inferences about unseen states in order to decide when to ask for help. Much past work has used Bayesian inference, which suffers tractability issues in complex environments. [^3] For the curious reader, [4] provides a thorough treatment. See also Section 2. In this paper, we instead assume that the mentor policy satisfies a Lipschitz-flavored condition we call “local generalization”: informally, if the mentor told us that action $a$ was safe in a nearby state, then $a$ is probably (but not certainly) also safe in the current state. Local generalization captures the intuition that one can transfer knowledge between similar states but the generalization becomes progressively less reliable as the states become less similar. Unlike Bayesian inference, local generalization only requires computing distances and is compatible with any state space which admits a distance metric.
@@ 1.2 Our assumptions
Unlike the standard online learning model, we assume that the agent does not observe payoffs. This is because the payoff in our model represents the chance of avoiding catastrophe on that time step. In the real world, one only observes whether catastrophe occurred, not its probability. [^4] One may be able to detect “close calls” in some cases, but observing the precise probability seems unrealistic.
1.3 Standard online learning
An overview of standard online learning is in order before discussing our results. In the standard model, the agent observes a state on each time step and must choose an action. An adversary then reveals the correct action, which results in some payoff for the agent. The goal is sublinear regret, i.e., the average regret per time step should go to 0 as $T\to\infty$. Figure 1 delineates the precise differences between the standard model and our model.
If the adversary’s choices are unconstrained, the problem is hopeless: if the adversary determines the correct action on each time step randomly and independently, the agent can do no better than random guessing. However, sublinear regret becomes possible if (1) the hypothesis class has finite Littlestone dimension [26], or (2) the hypothesis class has finite VC dimension [38] and the input is $\sigma$-smooth [^5] Informally, the adversarial chooses a distribution over states instead of a precise state. See Section 3. [17].
The goal of sublinear regret in online learning implicitly assumes catastrophe is impossible: the agent can make arbitrarily many (and arbitrarily costly) mistakes as long as the average regret per time step goes to 0. In contrast, we demand subconstant regret: the total probability of catastrophe should go to 0. Furthermore, standard online learning allows the agent to observe payoffs on every time step, while our agent only receives feedback on time steps with queries. However, access to a mentor (and local generalization) allows our agent to learn without trying actions directly, which is enough to offset all of the above disadvantages.
Objective Regret goal Feedback Mentor Local gen.
Standard model Sum of payoffs Sublinear Every time step No No
Our model Product of payoffs Subconstant Only from queries Yes Yes
Figure 1: Comparison between the standard online learning model and our model. On a technical level, sublinear vs subconstant regret is a more important distinction than the sum vs product of payoffs: for example, taking a logarithm can transform the product into a sum. However, the multiplicative objective and the goal of subconstant regret are tightly related, since the former’s interpretation as the chance of catastrophe is what gives rise to the latter: we want the total (not average) chance of catastrophe to go to 0.
@@ 1.4 Our results
Informally, we show that avoiding catastrophe with the help of a mentor and local generalization is no harder than online learning without catastrophic risk.
More precisely, we first show that in general, any algorithm with sublinear queries to the mentor has arbitrarily poor regret in the worst-case (Theorem 4.1). As a corollary, even when it is possible to avoid catastrophe with certainty, any algorithm either asks for help on a constant fraction of time steps or is nearly guaranteed to cause catastrophe. Unlike online learning where the general impossibility result is trivial (the agent might as well guess randomly given an unconstrained adversary), local generalization significantly limits the adversary’s power and necessitates a careful construction.
Next, we present a simple algorithm whose total regret and rate of querying the mentor both go to 0 as $T\to\infty$ when either (1) the mentor policy class has finite Littlestone dimension or (2) the mentor policy class has finite VC dimension and the state sequence is $\sigma$-smooth. Our algorithm can handle a multi-dimensional unbounded state space and does not need detailed access to the feature embedding, instead using two simple operations. It does need to know the mentor policy class, as is standard in online learning. We initially prove the theorem for binary actions (Theorem 5.2) and then reduce learning with (finitely) many actions to the binary action case (Theorem C.1).
Finally, Appendix D presents another algorithm with the same guarantees of subconstant regret and sublinear queries under different assumptions. Most importantly, this result applies to the class of thresholds on $[0,1]$, which is known to be hard for standard online learning (see Section 2). This shows that our problem and standard online learning do not have exactly the same difficulty, but rather our problem is no harder than online learning.
The rest of the paper is structured as follows. Section 2 discusses related work. Section 3 formally defines our model. Section 4 presents our negative result for general mentor policies. Section 5 presents our positive result for simple mentor policy classes. Proofs are deferred to the appendix.
2 Related work
@@
Despite the ubiquity of irreparable/irreversible costs in the real world, theoretical work on this topic remains limited. This may be due to the fundamental modeling question of how the agent should learn about novel states or actions without actually trying them.
The most common approach is to allow the agent to ask for help. This alone is insufficient, however: the agent must have some way to decide when to ask for help. A popular solution is to perform Bayesian inference on the world model, but this has two tricky requirements: (1) a prior distribution which contains the true world model (or an approximation), and (2) an environment where computing (or approximating) the posterior is tractable. A finite set of possible environments satisfies both conditions, but is unrealistic in many real-world scenarios. In contrast, our algorithm can handle an uncountable policy class and a continuous unbounded state space, which is crucial for many real-world scenarios in which one never sees the exact same state twice.
Bayesian inference combined with asking for help is studied by [8; 9; 25; 28]. We also mention [15; 29; 37], which utilize Bayesian inference in the context of safe (online) reinforcement learning without asking for help (and without regret bounds).
We are only aware of two papers which theoretically address irreversibility without Bayesian inference: [13] and [27]. The former proposes to sample trajectories and learn reversibility based on temporal consistency between states: intuitively, if $s_{1}$ always precedes $s_{2}$, we can infer that $s_{1}$ is unreachable from $s_{2}$. Although the paper theoretically grounds this intuition, there is no formal regret guarantee. The latter presents an algorithm which asks for help in the form of rollouts from the current state. However, the regret bound and number of rollouts are both linear in the worst case, due to the dependence on the $\gamma^{*}$ parameter which roughly captures how bad an irreversible action can be. In contrast, our algorithm achieves good regret even when actions are maximally bad.
@@
Online learning.
See [7] and Chapter 21 of [34] for introductions to online learning. A classical result states that sublinear regret is possible iff the hypothesis class has finite Littlestone dimension [26]. However, even some simple hypothesis classes have infinite Littlestone dimension, such as the class of thresholds on $[0,1]$ (Example 21.4 in [34]). Recently, [17] showed that if the adversary only chooses a distribution over states rather than the precise state, only the weaker assumption of finite VC dimension [38] is needed for sublinear regret. Specifically, they assume that each state is sampled from a distribution whose concentration is upper bounded by $\frac{1}{\sigma}$ times the uniform distribution. This framework is known as smoothed analysis.
Smoothed analysis.
Smoothed analysis was originally proposed by [36] to explain the fact that some algorithms almost always perform well in practice, despite the theoretical existence of pathological inputs. Informally, algorithms which perform well under smoothed analysis are guaranteed to perform well for nearly all – but not necessarily all – inputs.
Active learning.
Our assumption that the agent only receives feedback in response to queries falls under the umbrella of active learning [18]. This contrasts with passive learning, where the agent receives feedback automatically. Although ideas from active learning could be useful in our domain, we are not aware of any results from that literature which account for irreversible costs.
@@
Perhaps the best-studied example of a multiplicative objective is the Nash welfare function (also known as the Nash bargaining solution), defined as the geometric mean of agent utilities [23; 31]. Inspired by Nash welfare, [3] recently studied algorithms for maximizing the geometric mean of payoffs across time steps in the multi-armed bandit problem. However, the geometric mean of payoffs is much more forgiving of errors compared to the product of payoffs. Imagine that the agent obtains a payoff of $1-1/\sqrt{T}$ per time step: then $\lim_{T\to\infty}(1-1/\sqrt{T})^{T}=0$ but $\lim_{T\to\infty}((1-1/\sqrt{T})^{T})^{1/T}=\lim_{T\to\infty}(1-1/\sqrt{T})=1$. In our setting, it is crucial that the total chance of catastrophe (the product of payoffs) is small, not just the average chance of catastrophe per time step (the geometric mean of payoffs). This is closely related to the difference between sublinear and subconstant regret discussed in Section 1.3.
@@ 3 Model
States.
Let $\mathbb{N}$ refer to the set of strictly positive integers and let $T\in\mathbb{N}$ be the time horizon. Let $S\subseteq\mathbb{R}^{n}$ be the state space and let $\mathbf{s}=(s_{1},s_{2},\dots s_{T})\in S^{T}$ be the sequence of states. In the adversarial setting, each $s_{t}$ can have arbitrary dependence on the events of prior time steps. In the smoothed setting, the adversary only chooses the distribution $s_{t}$ from which $s_{t}$ is sampled. Formally, a distribution $\mathcal{D}$ over $S$ is $\sigma$-smooth if for any $X\subseteq S$, $\mathcal{D}(X)\leq\frac{1}{\sigma}U(X)$. (In the smoothed setting, we assume that $S$ supports a uniform distribution $U$ [^6] For example, $S$ having finite Lebesgue measure is sufficient. Note that this does not imply boundedness. Alternatively, $\sigma$-smoothness can be defined with respect to a different distribution, as long as the Radon-Nikodym derivative is uniformly bounded; see Definition 1 of [5].) If each $s_{t}$ is sampled from a $\sigma$-smooth $\mathcal{D}_{t}$, we say that $\mathbf{s}$ is $\sigma$-smooth. The sequence $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ can still be adaptive, i.e., the choice of $\mathcal{D}_{t}$ can depend on the events of prior time steps.
Actions.
Let $A$ be a finite set of actions. There also exists a special action $\hat{a}$ which corresponds to querying the mentor. For $k\in\mathbb{N}$, let $[k]=\{1,2,\dots,k\}$. On each time step $t\in[T]$, the agent must select an action $a_{t}\in\ A\cup\{\hat{a}\}$, which generates payoff $\mu(s_{t},a_{t})\in[0,1]$.
@@
The mentor is endowed with a (not necessarily optimal) policy $\pi^{m}:S\to A$. When action $\hat{a}$ is chosen, the mentor informs the agent of the action $\pi^{m}(s_{t})$ and the agent obtains payoff $\mu(s_{t},\pi^{m}(s_{t}))$. For brevity, let $\mu^{m}(s)=\mu(s,\pi^{m}(s))$. The agent never observes payoffs: the only way to learn about $\mu$ is by querying the mentor.
We would like an algorithm which becomes “self-sufficient” over time: the rate of querying the mentor should go to 0 as $T\to\infty$. Intuitively, this is equivalent to the cumulative number of queries being sublinear in $T$: $|Q_{T}|\in o(T)$, where $Q_{T}=\{t\in[T]:a_{t}=\hat{a}\}$. However, interpreted literally, $|Q_{T}|\in o(T)$ is not a rigorous statement. This is because $|Q_{T}|$ is not just a function of $T$: it can also depend on $\mathbf{s}$, $\mu$, $\pi^{m}$, and any randomness in the agent’s actions.
To formalize this, let $Q_{T}(\mathbf{s},\mu,\pi^{m})$ be the set of time steps with queries given $\mathbf{s},\mu$, and $\pi^{m}$. Then we say that $|Q_{T}|$ is sublinear in $T$, denoted $|Q_{T}|\in o(T)$, if there exists $g:\mathbb{N}\to\mathbb{N}$ such that $g(T)\in o(T)$ and $|Q_{T}(\mathbf{s},\mu,\pi^{m})|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{m}$ (and for all possible realizations of the agent’s randomization). However, we will mostly write $Q_{T}$ and $|Q_{T}|\in o(T)$ for brevity.
Local generalization.
We assume that the mentor policy permits “local generalization”. Formally, we assume there exists $L>0$ such that for all $s,s^{\prime}\in S$, $|\mu^{m}(s)-\mu(s,\pi^{m}(s^{\prime}))|\leq L||s-s^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance. Informally, if the agent is in state $s$, taking the mentor action from a nearby state $s^{\prime}$ is almost as good as taking the mentor action for $s$. In other words, it’s okay to use an action you learned from a nearby state.
This assumption might initially feel less intuitive than a standard Lipschitz continuity assumption of $|\mu(s,a)-\mu(s^{\prime},a)|\leq L||s-s^{\prime}||$. However, Lipschitz continuity is strictly stronger than local generalization when the mentor is optimal (Proposition E.1). The properties are incomparable in general, but Lipschitz continuity allows odd situations where the mentor provides an action which is only acceptable in that specific state. This sets the agent up for failure: even if such actions exist, a thoughtful mentor should avoid those actions when possible. See Figure 2 for an illustration.
[…]
Figure 2: The difference between local generalization and Lipschitz continuity. The x-axis corresponds to a one-dimensional state space, with $s_{1}$ and $s_{2}$ denoting two nearby states. The y-axis is the payoff. In all figures, $\pi^{m}(s_{1})=0$ and $\pi^{m}(s_{2})=1$. Left: What one might consider the “typical case” with a near-optimal mentor. The mentor takes action 0 in $s_{1}$, so $|\mu(s_{2},0)-\mu(s_{2},\pi^{m}(s_{2})|$ is small (local generalization) and $|\mu(s_{2},0)-\mu(s_{1},0)|$ is small (Lipschitz continiuty). Middle: Local generalization without Lipschitz continuity. Action 0 is much worse in $s_{2}$, but it is still nearly as good as the mentor’s action in $s_{2}$, which is also quite bad. We primarily care about the agent’s performance relative to the mentor. Right: Lipschitz continuity without local generalization. The payoff of each action is similar in nearby states. However, despite the mentor recommending action 0 in $s_{1}$, action 0 is much worse than the mentor’s action in $s_{2}$. The mentor is setting up the agent for failure by recommending action 0 in $s_{1}$.
@@
If $\mu(s_{t},a_{t})\in[0,1]$ is the chance of avoiding catastrophe on time step $t$ (conditioned on no prior catastrophe), then $\prod_{t=1}^{T}\mu(s_{t},a_{t})$ is the agent’s overall chance of avoiding catastrophe. [^7] Conditioning on no prior catastrophe means we do not need to assume that these probabilities are independent (and if catastrophe has already occurred, this time step does not matter). This is due to the chain rule of probability. For a fixed state sequence $\mathbf{s}$ and agent actions $\mathbf{a}=(a_{1},\dots,a_{T})$, the agent’s regret is
$R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})=\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
We will usually write $R_{T}=R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})$ for brevity. We will study the expected regret over any randomness in $\mathbf{s}$ and/or $\mathbf{a}$. We desire subconstant worst-case regret: the total (not average) expected regret should go to 0 for any $\mu$ and $\pi^{m}$. Formally, we want $\lim_{T\to\infty}\sup_{\mu,\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=0$.
VC and Littlestone dimensions.
VC dimension [38] and Littlestone dimension [26] are standard measures of learning difficulty which capture the ability of a hypothesis class (in our case, a policy class) to realize arbitrary combinations of labels (in our case, actions). We omit the precise dimensions since we only utilize these concept via existing results. See [34] for a comprehensive overview.